%) 11.2.19-T Question Help ! A researcher notes that, in a certain country, a disproportionate number of software millionaireB. Predict the number of software millionaire birthdays that will occur in a decade where the total number of births in this country is 35 million. ( round to 2 decimal places as needed)

C. Fit a simple linear regression model relating number (y) of software millionaire birthdays in a decade to number (x) of CEO birthdays. give the least  squares prediction equation. ( round to 2 decimal places as needed)

D. Predict the number of software millionaire birthdays that will occur in a decade where the total number of births in this country is 35 million
Answer #1

X Y XY XX YY 28.016 3 84.048 784.896 9 24.065 1 24.065 579.124 1 31.603 10 316.03 998.75 100 40.796 16 652.736 1664.31 256 38
a)
= α+ br ΣΥ *ΣΧΧ -ΣΧ* ΣΧΥ Η 44.0 και 6623.976 - 1963941584.985 • 8 -16,8986 * 6 * 6623.978 – 196.394 και 196.394 , ηΣΧΥ - ΣΧ*
b) x = 35
y = -16.90 +0.7435.0) û = 9.01

c)
X Y XY XX YY 2 3 6 4 9 2 1 2 4 1 22 10 220 484 100 35 16 560 1225 256 7 8 56 49 64 Ο 6 0 0 36 ΣΧ = 68.0,ΣΥ = 44.0 ΣΧΥ = 844.0
= α+ br ΣΥ * ΣXX – ΣΧ* ΣΧΥ ηΣΧ2 – (ΣΧ)2 44.0 και 1766.0 – 68.0 * 844.0 - 3.4012 * 6 * 1766.0 – 68.0 * 68.0 , ηΣΧΥ - ΣΧ *
d) x = 35
y = 3.40 +0.35(35.0) y = 15.55
Please hit thumps up if the answer helped you